# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v2

@@ 8 Container Density
Here, we delve into the concept of container density and its relevance to the properties of oriented graphs. We explore how regular exterior degrees interact with the overall density of containers and the implications for graph structure.
Having proved that the interior degrees double within a container and that is linked to container size decreasing, we are ready to tackle the overall Seymour Second Neighborhood Conjecture. First we need to establish some fundamental truths about containers and oriented graphs in these counterexamples to set up the rest of the paper and things that must be true.
@@
What we saw in Lemma 6.1 (Regular Interior Degrees) that the regularity of interior degrees was something we were able to prove while we were trying to prove that back arcs did not exist. This both made sense and was very helpful in that environment. We see here that the regularity of interior degrees is not restricted to the container itself, but also impacts the exterior arcs to the next container - forcing them to be regular as well.
@@
This lemma is a more extended form of Lemma 5.6 (Weak Decreasing Exteriors Lemma), but without the added dependencies. It gives us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named. Several examples and observations brought this to a conjecture, which brought it here. The question now becomes is there any way we can strengthen this result?
We demonstrated that the Decreasing Exteriors Lemma relies on two key factors: the elimination of back arcs and the doubling of interior degrees within each container. Furthermore, we established that the elimination of back arcs is directly dependent on the interior degrees doubling within those containers.
By proving the previous lemma, we not only confirm the doubling of interior degrees but also conclude two additional results: (1) back arcs are entirely eliminated, and (2) the exterior arcs exhibit the desired decreasing pattern.
@@
Proof.
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a node with the minimum degree in $G$, and assume that every node in $G$ satisfies the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and let $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$.
Our goal is to show that the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided that $(v_{0},x)\in G$ (i.e., $x$ is reachable from $v_{0}$).
By Lemma 8.3 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$ with respect to $x$, $|ext(v_{i-1},x)|$, is bounded above by:
$|ext(v_{i-1},x)|\leq d^{+}(v_{0})-i$
where $d^{+}(v_{0})$ is the out-degree of $v_{0}$. This gives a limit on the number of nodes that $x$ can connect to at distance $i+1$ from $v_{0}$.
Since the graph satisfies the Decreasing Sequence Property (DSP), we know that all nodes in the graph have a degree at least as large as $v_{0}$’s degree. Therefore, the degree of node $x$ satisfies:
$d(x)\geq d(v_{0})$
This means that node $x$ must have at least $i$ neighbors at distance $i+1$ from $v_{0}$.
The structure of the graph explicitly disallows back arcs, i.e., edges pointing to nodes at earlier distances from $v_{0}$. Thus, $x$ cannot connect to nodes at distances less than $i$ without violating the structure of the graph. This constraint eliminates the possibility of $x$ mapping to earlier-distance nodes.
Given that $x$ must connect to at least $i$ nodes, and since it cannot connect to earlier-distance nodes (due to the disallowance of back arcs), the only remaining option for $x$ is to map to interior nodes—nodes that are at distance $i$ or higher.
Thus, since $x$ must map to at least $i$ distinct nodes, and all these must be interior nodes (by the elimination of back arcs and exterior connections), we conclude that:
$|int(v_{i-1},x)|\geq i$
This shows that the interior degree of $x$ with respect to $v_{i-1}$ is at least $i$.
Therefore, we have shown that for a node $x$ at distance $i$ from $v_{0}$, the interior degree of $x$ with respect to $v_{i-1}$, $|int(v_{i-1},x)|$, is at least $i$, as required.
$|int(v_{i-1},x)|\geq i$
Now that the exterior has an upper bound on it’s size we can see how that impacts the interior degree. Because the degree of each node is still impacted by the minimum degree node, they still need to have their out-degrees go somewhere. Lemma 8.4 (Interior Degrees Lower Bound) compliments 8.3 (Strong Decreasing Exteriors Lemma) by saying thaw each decrease in the exterior degree will be matched by an increase in the interior degree.
@@ 9 Transitive Triangles and Seymour Diamonds
Here, $x$ points to both $y$ and $u$, and $y$ also points to $u$. The ”transitivity” comes from the fact that the direction of the edges follows a logical progression: $x\rightarrow y$ and $y\rightarrow u$ implies a pathway $x\rightarrow u$ via $y$. The concept of a transitive triangle is significant in understanding how the structure of the graph influences properties like out-degrees in $G^{2}$.
In a transitive triangle, $x$ has a direct path to both $y$ and $u$, and the existence of the edge $(y,u)$ implies that the out-degree of $x$ in $G^{2}$ could potentially increase because the paths from $x$ to nodes reachable from $y$ and $u$ are being counted. Understanding how transitive triangles are distributed within the graph can help in analyzing the neighborhood structures and their effects on the square graph, is critical for analyzing the Seymour’s Second Neighborhood Conjecture.
@@
The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $u_{1}$ and $u_{2}$, and these two out-neighbors both share the same out-neighbor, $y$. This is basically a four vertex oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and more importantly the Seymour’s Second Neighborhood Conjecture.
@@
This lemma highlights the relationship between a minimum degree node and the formation of transitive triangles, providing a foundation for discussing the implications related to the Seymour conjecture.
@@
1. There are forward arcs $u\to v$ and $u\to w$, 2. There are forward arcs $v\to y$ and $w\to y$, 3. No back arcs exist between containers.
This lemma effectively describes the conditions for forming a Seymour diamond and links it to the implications regarding the Seymour conjecture. Because we can extend the previous lemmas to encompass any nodes that have the DSP. This inclusion emphasizes that transitive triangles and Seymour diamonds have been integral to our analysis rather than overlooked.
@@
Understanding transitive triangles and Seymour diamonds enriches our analysis of oriented graphs, demonstrating the interplay between different graph structures. This exploration contributes to the overall argument and leads us to our main theorem.
@@ 10.1 Introduction
In this pivotal section, we present the main theorem derived from our analysis of oriented graphs. We articulate the theorem clearly, outlining its implications and the conditions under which it holds.
This paper has been a battle between the chosen minimum degree node and the Decreasing Sequence Property (DSP), with the former acting as a catalyst pushing things forward and the latter acting as a bottleneck, trying to hold things back. Along the way, we have proved several key results, most notably that:
  • Back arcs cannot exist,
  • Interior degrees double inside each container,
  • The interior degree doubling causes a decrease in exterior size.
This interaction between interior degree increases and exterior size decreases drives the structure of the graph. However, this degree increase is bounded by the bound $\frac{n-1}{2}$ because of the size of the containers, which provides a limit on the maximum average degree inside a container. Once the system reaches this maximum bound, the exterior degree of a parent container will double, causing a chain reaction where both interior and exterior degrees double, ultimately leading to the node’s overall degree doubling as well.
We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), the No Back Arcs Lemma from section 6, but slightly rephrased.
@@ 10.2 Theorem Statement
Corollary 10.1 (Degree Doubling by Container Density) captures the insight exploration our interior degree growth dynamics within shrinking containers in an oriented graph. The language highlights the inevitable ”collision” and ties it back to fundamental graph constraints. Hopefully, this framing gives a precise summary of why these graphs have a limited capacity to sustain both shrinking containers and rising interior degrees without contradictions.
The main theorem, algorithm, and corollary encapsulates the findings of our research, providing a crucial result that advances the understanding of oriented graphs. The proofs and implications of this theorem will be further explored in the following sections.
@@
The discovery of multiple degree-doubling nodes transforms the Seymour Second Neighborhood Conjecture from a question of existence to one of abundance. This result not only confirms the conjecture but reveals an inherent richness in the structure of oriented graphs. The presence of multiple nodes is tied to the emergence of complete graph structures as containers decrease in size, showcasing an element of mathematical beauty. Moreover, this has real-world applications: in network analysis, it suggests that influence or importance may not rest with a single node but instead within a clique or tightly connected group, offering new perspectives for studying collaborative dynamics and group influence.
The identification of regularity was key to understanding why multiple nodes’ degrees can double in certain graphs. This concept offers a powerful explanation for the phenomenon. The regularity is the key factor that helps explain why multiple nodes can have their degrees double, rather than the search for multiple nodes being inherently complex. The regularity is the central feature that ties everything together.
@@
Regularity implies that as nodes increase in distance from the starting node, their connectivity and interior degrees will follow a pattern. This pattern is key to understanding why degree doubling happens more naturally and why multiple nodes can experience it simultaneously, if the conditions allow.
@@ 10.3 Decreasing Sequence Property Path Algorithm
We now present a precise algorithm to identify nodes whose out-degree doubles in the square of an oriented graph, as per Seymour’s Second Neighborhood Conjecture (SSNC). This algorithm synthesizes results from the lemmas, theorem, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
@@ 11 Conclusion and Future Work
Solving a conjecture with such a long and debated history requires thorough validation. This paper represents a comprehensive effort to meet the rigorous standards of mathematical proof, supported by tools designed to aid understanding. Acknowledging that such results are subject to detailed scrutiny, this process is viewed as an essential step in establishing confidence and recognition within the mathematical community.
The GLOVER structure—short for Graph Level Order—represents a significant advancement in addressing long-standing open problems in graph theory, such as the Seymour conjecture, which remained unresolved for 34 years. While not all oriented graphs are inherently well-ordered, the GLOVER structure reveals that, within the context of the Seymour conjecture, oriented graphs become effectively well-ordered. This property arises because any deviation from such an ordering would lead to a degree-doubling effect, rendering the conjecture invalid.
By leveraging this induced well-ordering, the GLOVER structure enables tools like mathematical induction to be applied effectively, facilitating a solution to the Seymour conjecture. Beyond this specific case, the framework provides a foundation for exploring dual metrics in graph theory, such as prioritizing distance and degree, and opens the door to applications in clustering, ranking, and multi-metric optimization.
The GLOVER structure emerges as a groundbreaking framework for addressing problems that rely on dual metrics, offering a versatile tool for organizing and analyzing data across theoretical and practical contexts. While its utility was demonstrated here in the context of the Seymour conjecture—where the dual metrics of distance and degree played a pivotal role—it is important to recognize that the structure’s applications extend far beyond regularity or degree-based problems.
At its core, the GLOVER structure’s ability to prioritize and organize relationships based on dual metrics allows it to address challenges that require both exhaustive and hierarchical searches. Unlike local search methods traditionally constrained by single metrics, the GLOVER approach enables partitioning and clustering in ways that preserve relationships and provide deeper insights. For instance, applications in clustering, ranking, and network analysis—whether in social networks, employee rankings, or other multivariate datasets—can benefit from this structure.
Looking ahead, the GLOVER structure challenges us to investigate its applicability to other open problems in graph theory and computer science. The approach may uncover new insights in contexts where dual or multi-metric prioritization is key, highlighting the broader potential of this versatile framework.
Future research and development should explore these broader applications. Questions of complexity, scalability, and adaptability to non-graph-based datasets remain promising areas for investigation. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
@@ 14 Appendix 2: Applicability of the Graph Level Order
Traditionally, a total order on a set is a linear, one-dimensional comparison between elements. The key properties of total orders—reflexivity, transitivity, and antisymmetry—give us a clear, unambiguous ranking where every pair of elements is comparable. Graph Level Orders (GLOVER) extend this concept to n-dimensions within a leveled framework.
@@
GLOVER accomplishes this through the use of containers that not only arrange elements in sequence but also enable multiple levels of organization. Elements within each level maintain an internal order, while the containers themselves are comparable across ranked layers. This structure allows for sophisticated comparisons both within a single level and between levels, making GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
The examples provided in this paper illustrate concrete cases of GLOVER. However, it’s the abstract nature of GLOVER that sets it apart as a transformative mathematical concept. This structure allows us to conceptualize any set of ranked organized objects—not just graphs with regularly structured nodes, but also diverse entities, such as sets with multiple attributes or time-sequenced data. Within discrete mathematics, the ability to impose a multi-dimensional total order provides a richer understanding of relationships among elements. The order goes beyond pairwise comparisons to establish a multi-layered framework capable of supporting complex operations.
The true strength of GLOVER lies in its flexibility. It accommodates inherent ordering within single-level sets while also adapting to more intricate structures. By generalizing beyond domain-specific contexts like graph theory, GLOVER becomes a universal framework for imposing total order on any ranked structure, from simple ordered pairs to multi-dimensional data sets in fields like machine learning or clustering rankings.
The GLOVER structure also uniquely supports induction in graph-based proofs due to its total order, providing a new way to approach graph theory problems. Unlike traditional graph structures, which typically do not allow for induction, the total ordering of nodes and containers in GLOVER makes it possible to apply inductive reasoning effectively.
This expanded understanding of total orders paves the way for developing algorithms, proofs, and applications capable of handling complex, scalable data structures. GLOVER is not merely a form of ordering; it is a system of relationships that facilitates the manipulation, comparison, and analysis of multi-dimensional data through a clear and consistent layering.
The GLOVER framework, as initially defined, is tailored to address the specific requirements of the Seymour Second Neighborhood Conjecture. Its rigid structure-classifying arcs either as interior or exterior, and strictly prohibiting skipping levels-ensures clarity and sufficiency for analyzing second neighborhoods, However the framework can be extended to accomodate more general problems, including those with nested containers or more flexible neighbor relationships .
Nested containers introduce a hierarchy within individual containers, grouping nodes based on finer structural or relational properties. This extension enables the framework to address problems where additional levels of granularity within the same partition group are required.
Definition 14.1.
A nested container $C_{i,j}$ is a subset of a container $C_{i}$ defined by additional criteria beyond the qualifying metric that determined it into $C_{i}$. For example, nodes in $C_{i,j}$ might share specific adjacency patterns or connectivity property properties. Use Cases:
  • Clustering Problems: Nested Containers can represent hierarchal clustering within each distance group, allowing comparisons between subgroups.
  • Weighted Graphs: In graphs with edge weights, nested containers could group nodes based on weight thresholds within the same distance level.
Example:
Suppose $C_{2}$ contains nodes $x,y,z$ all at distance 2 from $v_{0}$. If $x$ and $y$ are connected but not $z$, $x$ and $y$ might form a nested container $C_{2,1}$, while $z$ forms $C_{2,2}$
In certain problems, arcs may not connect nodes in non-adjacent containers, effectively skipping levels in the distance hierarchy. While the SSNC prohibits such arcs to maintain a second-neighbor focus, other graph probems may allow or even require these connections.
  • Extension - Relax the exterior neighbors to include $C_{j})$ for $j>i+1$, introducing a new category of arcs, long-distance exterior arcs.
  • Use Cases -
    • Network Design - In communication networks, skipping levels might model direct long-range connections (for example satellite links).
    • Flow Problems - In transportation or flow networks, skipping levels could represent shortcuts or direct transfers.
  • Impact on GLOVER Properties - Allowing for skipping levels introducing new challenges
    • Order Maintenance - Total order within containers must be preserved despite the additional complexity of the long-distance connections.
    • New Arc Classifications - Introduce distinctions like short range exterior arcs (j = i + 1) and long range exterior arcs (j ¿ i + 1).
Exterior neighbors, as defined in the SSNC, connect nodes in adjacent containers. Expanding this definition allows for more flexible interactions between containers, broadening the applicability of the GLOVER framework.
  • Generalized Exterior Neighbors - for $x\in C_{i}$, an exterior neighbor $y$ can belong to any $j>i$. The specific range of $j$ will depend on the problem context.
  • Example Problem: Consider a graph where connections beyond second neighborhoods are of interest. Exterior neighbors could include nodes up to $k$-steps away, accomodating problems requiring $k$-neighborhood analysis.
While these extensions offer exciting generalizations, the SSNC’s specific focus on second neighbors necessittates a stricter interpretation of the GLOVER framework.
  • No skipping levels - Ensures that second neighbors are strictly defined nodes in $C_{2}$.
  • Rigid container definitions - prevents ambiguity in arc classifications or neighbor interactions.
  • Limited use of nested containers - The SSNC does not require additional granularity within distance groups, making nested containers optional for this specific problem.
Example 14.1.
Graph Theory - Seymour Second Neighborhood Conjecture: Graphs, which are normally seen as an unordered arrangement of nodes and edges, were able to be ordered with the GLOVER. We initially chose a minimum degree node and partitioned the nodes of the graph into containers based on their proximity to this node.
Partitioning nodes by distance allowed for definition of the interior degree and exterior degree which were very important to showing that a node’s overall degree would double.
Example 14.2.
Engineering - Dynamic Scheduling: In fields such as manufacturing, task scheduling often involves prioritizing independent tasks. With GLOVER, tasks are grouped into containers based on dependency levels and ordered according to priority. For example, preparatory tasks are prioritized in early containers, while finishing tasks are sequenced last, minimizing delays and optimizing task flow.
Partitioning groups tasks by dependency, while ordering between containers ensures essential tasks are completed first. Within each container, tasks can be further organized by resource requirements, promoting efficiency at each priority level.
Edges introduced at the container level to represent this dependency. Interior edges would show common dependencies between tasks in a common container, whereas exterior edges represent tasks of different priorities with common dependencies.
From this we can use graph theory concepts like centrality to better understand things like important tasks, important dependencies, how the containers (task tiers relate to one another) etc.
Example 14.3.
Engineering - Resource Management: In network resource allocation or energy grid management, nodes can be clustered by proximity to demand or criticality and ordered to prioritize high-need resources. This ordering ensures that essential nodes remain accessible, while resource usage remains balanced across the network.
Partitioning groups resources based on proximity to critical nodes, while ordering between containers directs allocations toward high-demand areas. Within each container, resources are accessed in order of current availability, optimizing responsiveness to fluctuating demands.
Example 14.4.
Engineering - Project Management: In complex projects involving multiple teams and dependencies, tasks can be clustered by project phase and ordered sequentially to prevent bottlenecks. Foundational tasks, such as design and planning, are prioritized early, while execution tasks follow, ensuring smooth progression through each phase.
Partitioning groups tasks by project phase, while ordering between containers facilitates the transition of tasks from one stage to the next. This ranked order remains stable despite potential project changes, enabling efficient coordination across large teams.
Example 14.5.
Cybersecurity - Threat Detection and Prioritization: In cybersecurity, ranked structures are often essential for organizing and prioritizing threats. Using an GLOVER framework, threats can be grouped into containers based on their urgency or potential impact. For example, critical threats with known vulnerabilities might be prioritized in early containers, while less immediate risks or less probable attack vectors are sequenced into subsequent layers. This ranked ordering helps ensure that the most pressing issues are handled first, thereby minimizing potential breaches or disruptions.
Partitioning into containers groups threats by their severity and probability, while ordering between containers facilitates a systematic approach to addressing risks from the highest priority down. Within each container, threats are further organized based on attributes like type, location in the network, or historical occurrence, allowing security teams to deploy resources effectively. This layered approach enhances the overall defense strategy by ensuring structured response and continuous monitoring across multiple risk levels.
Example 14.6.
Natural Language Processing - Ranked Text Classification: In NLP, ranked total order graphs can improve text classification by organizing data through multi-level categories. For instance, documents can be grouped into containers based on broad topics (e.g., news, entertainment, scientific research) with each container further ordered by subtopics like sports or politics under news, and biology or physics under science. This layered structure allows models to systematically refine classifications, progressing from general to specific categories, improving the accuracy and interpretability of the classification model.
Partitioning into containers groups text by broad categories, while ordering between containers prioritizes classification tasks from general to specific. Within each container, documents can be ordered by relevance, keywords, or thematic clustering, which supports tasks like topic modeling or intent recognition. This approach provides a clear and adaptable framework for complex, multi-layered categorization, enhancing the model’s ability to handle diverse and evolving datasets.
The versatility of the Graphs Level Order (GLOVER) structure hints at its applicability beyond this proof of the Seymour Second Neighborhood Conjecture. GLOVER has the potential to address several open problems and conjectures within discrete mathematics and graph theory, where ranked or layered ordering of elements can provide new insights or simplify complex relationships. Future work may reveal more connections between GLOVER and these longstanding challenges, offering new paths for research and exploration.

### v3

@@ 8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Exteriors Decreasing Lemma implies that containers will decrease in size. We need to see how that decrease in size will impact the degrees of the nodes.
@@
We proved Lemma 6.1 (Regular Interior Degrees) on our way to proving that back arcs did not exist. LABEL:refintdeg is an important property of the graph though that we are going to use to get other properties of the graph, like 8.1.
@@
Lemma (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma), to give us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named.
@@
Having already placed an upper bound on the exterior sizes and containers, it is only natural that the there is an opposite bound for the interiors. Once again the minimum degree node is the driving force behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
@@ 9 Transitive Triangles and Seymour Diamonds
Here, we have $x$ pointing to $y$ and $u$; and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles we might improperly count out second neighbors and thus not have a correct algorithm.
@@
The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is basically a four vertex oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and more importantly the Seymour’s Second Neighborhood Conjecture.
@@
Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm.
@@
1. There are forward arcs $x\to y$ and $y\to u$, 2. There are forward arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it. We didn’t know it because they they did not disrupt our algorithm and they are not a hindrance to our algorithm.
@@
Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs. Next we will blend everything together into our main results.
@@ 10.1 Introduction
Here is where we sum up all the lemmas we have derived from our analysis on oriented graphs. We will state the final algorithm and some final results.
So far, this battle has been a battle between a catalyst - the chosen minimum degree node, and the bottleneck - the Decreasing Sequence Property (DSP). We have proved some important things - Back arcs cannot exist; interior degrees must double inside each container, the exterior degrees decrease in size; interior degrees proportionally increase in size.
There is a limit though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
@@ 10.2 Theorem Statement
Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes. This was represented by the yin and yang effects early on. We see it full circle here. Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbor neighbors represented the size of the containers.
@@
The discovery of multiple degree-doubling nodes lifts the Seymour Second Neighborhood Conjecture from a mere question of existence to one in abundance. There were many who doubted that one even existed. Now there are many.
Second is the way we find it. It’s the presence of an oriented complete graph as these containers decrease in size. There’s an element of mathematical beauty there.
Then third we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person there may be a whole set of people, a tightly connected group.
Identifying regularity was key to understanding why multiple nodes’ degrees can double in certain graphs. The regularity is the key factor that helps explain why multiple nodes can have their degrees double. The regularity is the central feature that ties everything together.
@@ 10.3 Decreasing Sequence Property Path Algorithm
We will now give a pseudocode algorithm to identify the node whose out-degree doubles in the square of an oriented graph. This is the culmination of the lemmas, theorems, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
@@ 11 Conclusion and Future Work
The Seymour Second Neighborhood Conjecture started as a question in 1990. Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
With a solution now in hand, its time to begin looking forward. The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where noone follows back. The same question is of interest, is there anyone whose degree doubles, except we know the answer is yes. We just need to implement the algorithm to find the person
We could have the graph represent epidemiology. Here the nodes be infected people, and the arcs would be one way infections. The question of interest would be can we find the person who is a super-spreader.
Secondly, we introduced a data structure, Graph Level Order-GLOVER. This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for GLOVER - the well ordering and the dual metrics of distance and degrees.
Future research and development should explore questions of complexity, scalability. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
@@ 14 Appendix 2: Applicability of the Graph Level Order
A total order is a linear ranking. There are no containers or buckets to make extra comparisons. Everything is sorted. That may not be what we desite though. Suppose we have two metrics, like age and income. Those two metrics cannot be compared on the same scale. They’d need to be grouped by one metric and compared by the other.
@@
The Graph Level Order (GLOVER) is able to do these multi level comparisons. GLOVER does this by the usage of containers. These containers arrange elements in sequence. Elements within each level maintain an internal order. The containers themselves are comparable across ranked layers. This structure allows for multi-level comparisons. These can happen both within a single level and between levels. This makes GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
Just as the SSNC was proved with the GLOVER data structure, there are more proofs and algorithms that this method is capable of handling.